# Apparate: Rethinking Early Exits to Tame Latency-Throughput Tensions in ML Serving

### v1

@@ 2.3 Challenges
Figure 6: Optimal EE configurations change frequently. Streaming workloads were divided into chunks of 64 requests. Dot presence show a ramp that was part of the optimal config for a chunk, while transparencies indicate threshold values (opaque is higher).
@@
Table 1: Thresholds need frequent tuning to avoid accuracy loss. Continual tuning kicks in when chunk accuracy $<$ 99%. Results list avg accuracy (median latency wins).
@@
C2: Frequent and costly adaptation. As shown in Figure 6, the evolving nature of workloads for interactive applications [51, 35] brings frequent changes in the best EE configuration at any time, i.e., the set of active ramps (and their thresholds) that maximize latency savings without sacrificing response accuracy or exceeding available memory. Unfortunately, the large body of EE literature is unaccompanied by any policy for tuning ramps and thresholds during serving. Instead, proposed EE models are equipped with the max number of ramps, and mandate users to perform one-time tuning of thresholds. Such tuning is non-trivial and fails to cope with workload dynamism. For example, Table 1 shows how one-time tuning on initial or sampled data brings 8.3-14.5% drops in accuracy relative to continual tuning. Worse, the space of configurations is untenably large, with many ramp options (i.e., at any layer, with any computation) and a continuous space of possible threshold values for each.
@@
C3: Lack of accuracy feedback. EE ramp decisions are ultimately confidence-driven and may result in accuracy degradations (as shown above). In production scenarios, serving optimizations that deliver accuracy reductions of more than 1-2% are generally considered unacceptable [14]. Yet, despite this strict constraint, once deployed, EE models do not provide any indication of accuracy drops; indeed, when an exit is taken, the corresponding input does not pass through the remaining model layers, and the original (non-EE) model’s prediction is never revealed. Thus, with early exiting, we lack mechanisms to determine when accuracy degradations are arising and EE tuning is required.
C4: Incompatibility with batching. Lastly, although EE decisions do not directly conflict with queuing decisions in serving platforms, combining exiting and batching presents practical challenges. In particular, as inputs exit at ramps, batch sizes for already scheduled inference tasks naturally shrink, leading to resource underutilization for the rest of the model’s execution. Batch reforming [4] could help, but adds undue latency from added CPU-GPU data transfers; recent proposals [38, 29] aim to address these communication overheads but fail to generalize across hardware configurations. [^1] Solving this challenge directly is outside the scope of this paper. We note that this would bring latency and compute wins from EEs, but would forego accuracy feedback (and thus, adaptation).
3 Design
@@
Apparate is an end-to-end system that automatically integrates early exits into models and manages their operation throughout the inference process. Its overarching goal is to optimize per-request latencies while adhering to tight accuracy constraints and throughput goals. Our key insight is in rethinking the way that EEs are configured and the benefits they are expected to deliver. In particular, rather than using EEs in the traditional way – where inputs exit model inference to provide both latency and computational benefits – Apparate focuses solely on latency savings by allowing results to exit, with inputs still running to completion. Foregoing true exiting (and thus, compute savings) eliminates batch size changes during inference (C4), while also granting Apparate with direct and continual feedback on EE accuracy (C3). This feedback, in turn, provides the requisite signals for Apparate’s strategy to continually adapt EE configurations to maximize latency savings while catering to resource constraints and workload dynamics (C1, C2).
Figure 7: System architecture.
@@
Figure 7 overviews Apparate’s workflow, which runs atop existing serving platforms. Users register inference jobs as normal 1, providing models and SLOs. In addition, Apparate introduces two parameters: (1) ramp aggression, which (along with compute restrictions) bounds the number of active ramps in terms of % impact on worst-case latency (and throughput), and (2) accuracy constraint which indicates how much (if any) accuracy loss is acceptable relative to running the submitted model on all inputs without exiting. With these inputs, Apparate’s controller begins by configuring the provided model with EEs 2, performing a graph-level assessment to determine suitable positions for ramps, and training those ramps on bootstrap data (§3.1). The resulting model is passed to the serving platform for deployment 3, after which Apparate shifts to management mode. In this phase, as requests arrive and inference tasks are scheduled, Apparate’s controller gathers real-time feedback on both the utility of each ramp (overheads vs. latency savings) and achieved accuracies (relative to the original model) 4. This information is continually used to adapt the EE configuration 7 at different time scales: rapid threshold tuning for accuracy preservation (§3.2) 5, and less frequent ramp adjustments for latency optimization (§3.3) 6.
@@
Note that Apparate’s controller runs on a CPU, with GPUs streaming per-ramp/batch profiling information in a non-blocking fashion. This is possible since inputs pass to the end of models with Apparate, irrespective of exiting decisions. Further, tasks associated with model handling and serving are handled by the underlying serving platform, e.g., loading from disk, queuing.
3.1 Preparing Models with Early Exits
@@
Upon job registration with a generic DNN, Apparate’s initial task is to automatically prepare that model to leverage EEs, without requiring any developer effort. This phase repeats any time the submitted model changes, e.g., continual retraining to cope with data drift [13, 51, 35].
@@
Ramp locations. Apparate accepts a model in the ONNX format, a widely used IR that represents the computation as a directed acyclic graph [6]. Once ingested, Apparate must first identify candidate layers for ramp addition. The goal is to maximize ramp coverage across the model (to provide more configuration options for Apparate’s runtime management), while avoiding ramps that are unlikely to be fruitful (but add complexity adaptation decisions). To balance these aspects for diverse models, Apparate marks feasible ramp locations as those where operators are cut vertices, i.e., a vertex whose removal would disconnect a graph into two or more disjoint sub-graphs. In other words, no edge can start before a ramp and re-enter the model’s computation after the ramp.
@@
Figure 8: Apparate only injects ramps that make full use of available data flows at that part of the model.
@@
The idea is that such ramps take advantage of all available data outputs from the original model’s processing to that point, boosting their chance at accurate predictions. As an example, consider families like ResNet or BERT which enable deep models by stitching together series of residual blocks, i.e., ResNet blocks for convolutions, or BERT encoders that each embed multi-head attention and feed-forward network residual blocks. To avoid performance degradations late in the model, the output of each block is ultimately a combination of its processing results and its input. In such scenarios, Apparate injects ramps between blocks, but not within each block to avoid ramps making decisions on partial data, i.e., ignoring block inputs. In contrast, for VGG models, ramps are feasible at all layers since their intermediates represent the full extent of data flow throughout the model. Figure 8 depicts these examples.
@@
Overall, this strategy results in 9.2-68.4% of layers having ramps for the models in our corpus, which we empirically observe is sufficient to adapt to dynamic workloads (§5.2). However, we note that Apparate can directly support any other ramp configuration strategy, and offers a simple API for developers to express ramp policies or restrictions. Moreover, for each feasible ramp, Apparate’s runtime monitoring (§3.2-3.3) ultimately monitors its accuracy and utility (and thus, if it is ever activated) for the workload at hand.
@@
Apparate opts for the shallowest ramps that can transform the intermediates at any layer into a final model prediction. Specifically, ramps comprise the model’s final fully-connected (fc) layer, prepended with a lightweight pooling operation that reduces the dimensionality of intermediates to ensure compatability with the fc layer. This manifests differently for various model types. For instance, for vision models like ResNet, pooling is simply the model’s penultimate layer. In contrast, for BERT, only the basic operator is drawn from the BERT pooler module, i.e., extracting the hidden state corresponding to the first token [19]. Note that, for all models, the input width of the fc layer is modified to match that of the intermediates at each ramp location; the output remains unchanged to preserve result formats.
@@
Figure 9: More lightweight ramps boost EE savings. Results compare Apparate’s default ramps (many ramps) with versions that use fewer, more expensive ramps, i.e., adding convolutional layers for CV, or more fully-connected/pooling layers for NLP.
@@
Figure 9 evaluates this methodology by comparing Apparate’s ramps with two, more expensive alternatives. With ResNet, to mimic model operations following each ramp, we add 1-2 convolution layers prior to pooling. For BERT, we consider two approaches: (1) add two fc layers after pooling, each with reduced width to shrink inputs to the final fc, and (2) inspired by DeeBERT [57], replacing the simple pooling operator with the entire BERT pooler block (which also includes dense linear and activation layers) and adding a dropout after pooling as in the original model. In all cases, the number of ramps is subject to the same ramp budget (i.e., Apparate’s default uses the most ramps), ramps are uniformly spaced across feasible positions in each model, and thresholds are optimally selected as in §2.2.
@@
As shown, and in line with reflections from prior hand-designed EE models [53], we observe that the additional computation has minimal effect on ramp effectiveness. For example, median latencies are 1.3-1.8$\times$ and 1.9-5.4$\times$ larger with Apparate’s default ramps than the more complex alternatives for CV and NLP. Nonetheless, to showcase Apparate’s generality, we consider other ramp styles in §5.4.
@@
Training ramps and deploying models. To determine the appropriate weights for each ramp, Apparate relies on an initial dataset that can either be user-provided or auto-generated by running the original model on historical data. Regardless, during training, Apparate freezes the original model weights to ensure that non-EE behavior and feedback for tuning EEs is unchanged from the user’s original intentions. In addition, Apparate enforces that all inputs are used to train all ramps, i.e., exiting is prohibited during training. This ensures that ramps are trained independently of the presence (or behavior) of any upstream ramps, which is crucial since the set of active ramps can vary at runtime. Further, such independence and the model freezing enable loss calculations to be backwards propagated in parallel across ramps, rapidly speeding up training despite Apparate’s use of many lightweight ramps. §4 details the training process.
@@
For initial deployment, Apparate evenly spaces the max number of allowable ramps (based on the budget and GPU resources) across the model. To avoid accuracy dips due to discrepancies between training data and the current workload, each ramp begins with a threshold of 0, i.e., no exiting. The updated model definition (with enabled ramps) is passed to the serving platform which operates as normal, e.g., profiling expected model runtimes for different batch sizes [22].
3.2 Accuracy-Aware Threshold Tuning
@@
To enable threshold tuning, as requests pass through a model, Apparate’s controller continually records exiting information at each active ramp, as well as the final result that the original model predicts. More precisely, for each ramp, Apparate records the result (e.g., classification label) with the lowest error rate, i.e., the highest-confidence result for the ramp, even if the error exceeds the ramp’s threshold (precluding exiting). Importantly, since inputs always pass fully through models with Apparate, this information is recorded for all inputs at each active ramp, irrespective of upstream exiting decisions. This is paramount since the information serves not only as a signal for when to tune thresholds, but also provides guidance for how to do so.
@@
Evaluating threshold configurations. Threshold tuning requires insight into how any alterations to active ramp thresholds would affect overall model exiting behavior (and accuracies). The aforementioned per-request, per-ramp monitoring information grants this visibility, enabling Apparate to rapidly evaluate any threshold values across active ramps without requiring additional inference, and while accounting for inter-ramp dependencies. In particular, to evaluate new threshold values, Apparate simply identifies the earliest ramp whose top prediction now has an error rate below its threshold. Comparing these results with those of the original model indicates the achieved accuracy for the new configuration; latency wins for these exit patterns are computed using the one-time profiling data described in §3.3.
@@
Greedy search. The goal of tuning is to identify a new set of thresholds that maximize latency savings while adhering to accuracy constraints for the last window of data. The challenge is that the space of thresholds to consider is massive, precluding a grid search (especially given how frequently adaptation is needed - §2.3). Indeed, even with discretized threshold values in [0, 1] with a step size of $S$, computation costs are $O(C\times(\frac{1}{S})^{R})$, where $R$ is the number of active ramps, and $C$ is the cost to evaluate a given configuration.
@@
Figure 10: Threshold tuning example with two active ramps for ResNet50 and a random video. Configurations within the boundary have $<$1% accuracy loss; cell values list latency wins. Arrows show the path taken by Apparate’s hill climbing algorithm (without fine-grained step changes).
@@
Instead, Apparate employs a greedy heuristic (Algorithm 1 in §A) that leverages a fundamental property of EEs when evaluated against an original model: higher thresholds result in monotonic decreases in accuracy and monotonic increases in latency savings. This prunes the space of threshold values to consider by providing a clear boundary in the $R$-dimensional space that separates configurations that are sufficiently accurate from those that are not. Additionally, for accurate configurations, maximum latency savings must fall on that boundary. Figure 10 illustrates this.
@@
These properties inform Apparate’s hill climbing strategy [47] for threshold tuning. Starting with threshold values of 0 for each active ramp, and a step size of 0.1, threshold tuning runs in a series of (incremental) exploration rounds. In each round, we increase the threshold of each ramp in isolation (leaving the others unchanged), and evaluate the achieved accuracy and latency savings as described above. Apparate then chooses the single ramp threshold change that delivered the largest additional latency savings per unit of additional accuracy loss. This process repeats until no ramp’s threshold can be increased without an accuracy violation.
@@
To enhance this process, Apparate follows a multiplicative increase, multiplicative decrease policy on step sizes to balance search speed and granularity. Specifically, each time a step increase results in an accuracy violation, Apparate halves that ramp’s step size for subsequent rounds to hone in on the boundary at fine granularity; step sizes are lower-bounded at 0.01. Conversely, selection of a ramp for threshold alteration suggests a potentially promising path of exploration; in this case, for a speedup, Apparate doubles that ramp’s step size for the following round.
@@
Overall, as shown in Figure 11, Apparate’s threshold tuning algorithm runs up to 3 orders of magnitude faster than a pure grid search (11.9ms vs. 3.0s on average). Note that these results maximally parallelize grid search across a 30-core machine. Further, selected threshold values achieve within 0-3.8% of the latency savings of the optimal configurations.
@@
Figure 11: Apparate’s tuning vs. optimal tuning on runtime and latency of selected configurations. Bars list medians across all model-workload pairs, with error bars for min-max.
3.3 Latency-Focused Ramp Adjustments
The set of active ramps ultimately dictates where inputs can exit, and thus provides bounds on potential latency savings. Unlike threshold tuning which runs reactively (since accuracy is a constraint) and uses only recent profiling data to evaluate new configurations, ramp adjustment is strictly an optimization (for latency savings), and requires deployment to evaluate the impact of any new ramp. Thus, Apparate’s ramp tuning runs periodically (every 128 samples by default) and conservatively alters the set of active ramps to incrementally converge on high-performing configurations.
Evaluating active ramps. In each round, Apparate’s controller starts by computing a utility score for each active ramp that evaluates its overall impact on workload latency. To do so, Apparate couples per-ramp exit rates (based on profiling data from threshold tuning in §3.2) with two additional inputs that are collected once per model during bootstrapping: (1) the latency overhead per ramp, and (2) a layer-wise breakdown of time spent during model inference (for different batch sizes [22]). The latter is necessary since different models can exhibit wildly different latency characteristics that govern the impact of any exits. For instance, latency of CV models is often heavily skewed towards early layers given the high dimensionality of the input data [40], while NLP transformers exhibit more consistent latency values across coding blocks. Note that, in the case of distributed serving, latency breakdowns are updated to account for network delays between serving nodes.
Using these inputs, Apparate defines the utility of ramp $R$ as savings - overheads, where savings represents the sum of raw latency that exiting inputs avoided by using ramp $R$, and overheads is the sum of raw latency that $R$ added to inputs that it was unable to exit. A negative (positive) utility value means a ramp is causing more (less) harm than benefit to overall latencies for the current workload.

### v2

@@ 2.3. Challenges
Figure 5. Optimal EE configurations change frequently. Workloads use 64-request chunks. Dot presence shows a ramp that was part of the optimal config for a chunk, while transparencies indicate threshold values (opaque is higher).
@@
Table 1. Thresholds need frequent tuning to avoid accuracy loss. Continual tuning kicks in when chunk accuracy $<$ 99%. Results list avg accuracy (median latency wins).
@@
C2: Frequent and costly adaptation. As shown in Figure 5, the evolving nature of workloads for interactive applications brings frequent changes in the best EE configuration at any time, i.e., the set of active ramps (and their thresholds) that maximize latency savings without sacrificing response accuracy. Unfortunately, the large body of EE literature is unaccompanied by any policy for tuning ramps and thresholds during serving. Instead, proposed EE models are equipped with the max number of ramps, and mandate users to perform one-time tuning of thresholds. Such tuning is non-trivial and fails to cope with workload dynamism. For example, Table 1 shows how one-time tuning on sampled data brings 8.3–14.5% drops in accuracy relative to continual tuning. Worse, the space of configurations is untenably large, with many ramp options (i.e., at any layer, with any computation) and a continuous space of possible threshold values for each.
@@
C3: Lack of accuracy feedback. EE ramp decisions are ultimately confidence-driven and may result in accuracy degradations (as shown above). In production scenarios, serving optimizations that deliver accuracy reductions $>$1–2% are generally considered unacceptable (Bolukbasi et al., 2017). Yet, once deployed, EE models do not provide any indication of accuracy drops; indeed, when an exit is taken, the corresponding input does not pass through the remaining model layers, and the original (non-EE) model’s prediction is never revealed. Thus, with early exiting, we lack mechanisms to determine when accuracy degradations are arising and EE tuning is required.
3. Design
@@
Apparate is an end-to-end system that automatically integrates early exits into models and manages their operation throughout the inference process. Its overarching goal is to optimize per-request latencies while adhering to tight accuracy constraints and throughput goals. Our key insight is in rethinking the way that EEs are configured and the benefits they are expected to deliver. In particular, rather than using EEs in the traditional way – where inputs exit model inference to provide both latency and computational benefits – Apparate focuses solely on latency savings by allowing results to exit, with inputs still running to completion. Foregoing true exiting (and thus, compute savings) grants Apparate with direct and continual feedback on EE accuracy (C3). This feedback provides the requisite signals for Apparate to continually adapt EE configurations to maximize latency savings while catering to resource constraints and workload dynamics (C1, C2).
Apparate’s design represents a departure from the typical expected utility of EEs (i.e., compute savings) that has been fraught with practical challenges. Instead, Apparate demonstrates an alternate avenue for benefits that EEs can bring (i.e., latency reductions), while remaining compatible with other compute efficiency optimizations that have had substantial practical traction. For instance, by foregoing true exiting, Apparate can run alongside request batching (Iyer et al., 2024). Further, Apparate supports diverse model architectures, including those that have been compressed for efficiency (§4.2). We note that the redundant computations in Apparate match the work that vanilla serving perform by executing all model layers.
Figure 6. System architecture.
@@
Figure 6 overviews Apparate’s workflow, which runs atop existing serving platforms. Users register inference jobs as normal 1, providing models and SLOs without needing any awareness of or expertise about EEs. In addition, Apparate introduces two parameters: (1) ramp aggression, which bounds the number of active ramps in terms of % impact on worst-case latency (and throughput), and (2) accuracy constraint which indicates how much (if any) accuracy loss is acceptable relative to running the submitted model on all inputs without exiting. Apparate’s controller begins by configuring the provided model with EEs 2, performing a graph assessment to determine suitable positions for ramps, and training those ramps on bootstrap data (§3.1). The resulting model is passed to the serving platform for deployment 3, after which Apparate shifts to management mode. In this phase, as requests arrive and inference is scheduled, Apparate’s controller gathers real-time feedback on the utility of each ramp (overheads vs. latency savings) and achieved accuracies (relative to the original model) 4. This data is used to continually adapt the EE configuration 7 at different time scales: rapid threshold tuning for accuracy preservation (§3.2) 5, and less frequent ramp adjustments for latency optimization (§3.3) 6.
This decoupling of EE configuration adaptation into two tuning control loops is a key design decision that Apparate uses to manage the untenably large search space of ramp-threshold combinations (C1) without substantial loss in EE efficacy. Specifically, Apparate chooses to use frequent threshold tuning to preserve accuracy because it provides a finer-grained knob for walking the accuracy-latency tradeoff, i.e., thresholds are continuous, whereas ramp locations are inherently discretized. Thresholds also provide a mechanism to control ramp location; at the extreme, thresholds for any active ramps can be tuned to preclude exiting. Regardless, to limit foregone wins from infrequent ramp tuning, Apparate opts to employ many lightweight ramps (§3.1): even if an optimal ramp is not present yet, a nearby ramp is likely active and can provide much of the same wins.
@@
Implementation details. Apparate is implemented as a layer atop existing serving platforms (currently three (Olston et al., 2017; Gujarati et al., 2020; hf_, 2024), though its techniques generalize to others), and comprises $\sim$7500 lines of Python code for EE preparation (§3.1) and management (§3.2-3.3). Apparate runs a separate controller per model replica (as decided by the serving platform) on a CPU, with GPUs streaming per-ramp/batch profiling information in a non-blocking fashion. This is possible since inputs pass to the end of models with Apparate, irrespective of exiting decisions. Tasks associated with model handling and serving are handled by the underlying serving platform, e.g., loading from disk, queuing, and batching.
3.1. Preparing Models with Early Exits
@@
Upon job registration with any DNN, Apparate’s initial task is to automatically prepare the model to leverage EEs without developer effort. This phase repeats any time the submitted model changes, e.g., continual retraining (Bhardwaj et al., 2022; Suprem et al., 2020; Khani et al., 2023). Note that, in the event that a developer provides an EE model, Apparate can forego any training and instead immediately begin managing its exit configurations (§3.23.3).
@@
Ramp locations. Apparate accepts a model in the ONNX format, a widely used IR that represents the computation as a directed acyclic graph (onn, 2024b). Once ingested, Apparate must first identify candidate layers for ramp addition. The goal is to maximize ramp coverage across the model (to provide more configuration options for Apparate’s runtime management), while avoiding ramps that are unlikely to be fruitful (but add complexity to adaptation decisions). To balance these aspects for diverse models, Apparate marks feasible ramp locations as those where operators are cut vertices, i.e., a vertex whose removal would disconnect a graph into two or more disjoint sub-graphs. In other words, no edge can start before a ramp and re-enter the model’s computation after the ramp.
@@
Figure 7. Apparate only injects ramps that make full use of available data flows at that part of the model.
@@
The idea is that such ramps take advantage of all available data outputs from the original model’s processing to that point, boosting their chance at accurate predictions. As an example, consider families like ResNet or BERT which enable deep models by stitching together series of residual blocks, i.e., ResNet blocks for convolutions, or BERT encoders that each embed multi-head attention and feed-forward network residual blocks. To avoid performance degradations late in the model, the output of each block is ultimately a combination of its processing results and its input. In such scenarios, Apparate injects ramps between blocks, but not within each block to avoid ramps making decisions on partial data, i.e., ignoring block inputs. Ramps are similarly injected between trasnformer blocks of generative models, but only for decoding phases (as input tokens must be fully processed for generation). In contrast, for VGG models, ramps are feasible at all layers since their intermediates represent the full extent of data flow throughout the model. Figure 7 depicts examples.
@@
Overall, this strategy results in 9.2–68.4% of layers having ramps for the models in our corpus, which we empirically observe is sufficient to adapt to dynamic workloads (§4.2). However, we note that Apparate can directly support any other ramp configuration strategy, and offers a simple API for developers to express ramp policies or restrictions.
@@
Apparate opts for the shallowest ramps that can transform the intermediates at any layer into a final model prediction. Specifically, ramps comprise the model’s final fully-connected (fc) layer, prepended with a lightweight pooling operation that reduces the dimensionality of intermediates to ensure compatability with the fc layer. This manifests differently for various model types. For instance, for vision models like ResNet, pooling is simply the model’s penultimate layer. Similarly, for generative LLMs, ramps can simply use the final decoder head to transform intermediate hidden states. In contrast, for BERT, only the basic operator is drawn from the BERT pooler module, i.e., extracting the hidden state corresponding to the first token (Devlin et al., 2018). For all models, the input width of the fc layer is modified to match the intermediates at each ramp location; the output remains unchanged to preserve result formats.
@@
Figure 8. More lightweight ramps boost EE savings. Results compare Apparate’s default ‘many ramps’ with versions that use fewer, more expensive ramps.
@@
Figure 8 evaluates this methodology by comparing with two, more expensive alternatives. With ResNet, to mimic model operations following each ramp, we add 1–2 convolution layers prior to pooling. For BERT, we consider two approaches: (1) add two fc layers after pooling, each with reduced width to shrink inputs to the final fc, and (2) following DeeBERT (Xin et al., 2020), replace the simple pooling operator with the entire BERT pooler block and add a dropout as in the original model. In all cases, the number of ramps is subject to the same budget (i.e., Apparate’s default uses the most ramps), ramps are uniformly spaced across feasible positions in each model, and thresholds are optimally selected as in §2.2.
@@
We observe that the added compute has minimal effect on ramp efficacy. For example, median latencies are 1.3–1.8$\times$ and 1.9–5.4$\times$ smaller with Apparate’s default ramps than the complex alternatives for CV and NLP. Nonetheless, to show Apparate’s generality, we consider other ramp styles in §4.5.
@@
Training ramps and deploying models. To determine the appropriate weights for each ramp, Apparate can automatically label a dataset that is either developer-provided or collected online (running the vanilla model on live inputs). Automatic labeling is feasible since ramps aim to mimic the submitted model’s behavior (not ground truth), so the submitted model’s outputs can directly serve as labels. Regardless, during training, Apparate freezes the original model weights to ensure that non-EE behavior and feedback for tuning EEs is unchanged from the user’s original intentions. Since its ramps are lightweight (single fully-connected layers) and comprise only 0.01–3.50% of our models’ parameters, the FLOPs required for Apparate’s ramp training is significantly lower than whole-model pre-training or even fine-tuning. In cases where existing (final) layers can be used as ramps, e.g. in generative scenarios, Apparate eschews training and directly reuse the final layer for each ramp. In addition, Apparate enforces that all inputs are used to train all ramps, i.e., exiting is prohibited during training. This ensures that ramps are trained independently of the presence (or behavior) of any upstream ramps, which is crucial since the set of active ramps can vary at runtime. Further, such independence and model freezing enable loss calculations to be backwards propagated in parallel across ramps, rapidly speeding up training despite the many lightweight ramps. As a result, ramp training only takes on the order of a few minutes for our models using a single A6000 GPU. Apparate uses the first 10% of each dataset for training and validation (following a 1:9 split).
@@
For initial deployment, Apparate evenly spaces the max number of allowable ramps across the model. To avoid accuracy dips due to discrepancies between training data and the current workload, each ramp begins with a threshold of 0, i.e., no exiting. The updated model definition (with enabled ramps) is passed to the serving platform which runs normally.
3.2. Accuracy-Aware Threshold Tuning
@@
To enable threshold tuning, as requests pass through a model, Apparate continually records exiting information at each active ramp, as well as the final result that the original model predicts. More precisely, Apparate records the highest-confidence result for each ramp, even if the error exceeds the ramp’s threshold (precluding exiting). Importantly, since inputs always pass fully through models with Apparate, this information is recorded for all inputs at each active ramp, irrespective of upstream exiting decisions. This is paramount since the information serves not only signals when to tune thresholds, but also provides guidance for how to do so.
@@
Evaluating threshold configurations. Threshold tuning needs insight into how any alterations to active ramp thresholds would affect model exiting behavior (and accuracies). By observing per-request behavior only at active ramps, Apparate can rapidly evaluate any threshold configuration without additional inference, and while accounting for inter-ramp dependencies. In particular, to evaluate new threshold values, Apparate simply identifies the earliest ramp whose top prediction now has an error rate below its threshold. Comparing these results with those of the original model indicates the achieved accuracy for the new configuration; latency wins are computed using the one-time profiling data described in §3.3.
@@
Greedy search. The goal of tuning is to identify a new set of thresholds that maximizes latency savings while adhering to accuracy constraints for the last window of data. The challenge is that the space of thresholds to consider is massive, precluding a grid search (especially given how frequently adaptation is needed - §2.3). Indeed, even with discretized threshold values in [0, 1] with a step size of $S$, computation costs are $O(C\times(\frac{1}{S})^{R})$, where $R$ is the number of active ramps, and $C$ is the cost to evaluate a given configuration.
@@
Figure 9. Threshold tuning example with two active ramps for ResNet50 and a random video. Configurations within the boundary have $<$1% accuracy loss; cell values list latency wins. Arrows show the path taken by Apparate’s hill climbing algorithm.
@@
Instead, Apparate employs a greedy heuristic that leverages a fundamental property of EEs when evaluated against an original model: higher thresholds result in monotonic decreases in accuracy and monotonic increases in latency savings. This prunes the space of threshold values to consider by providing a clear boundary in the $R$-dimensional space that separates configurations that are sufficiently accurate from those that are not. Additionally, for accurate configurations, maximum latency savings must fall on that boundary. Figure 9 illustrates this.
@@
These properties inform Apparate’s hill climbing strategy (Selman and Gomes, 2006) for threshold tuning. Starting with threshold values of 0 for each active ramp, and a step size of 0.1, threshold tuning runs in a series of (incremental) exploration rounds. In each round, we increase the threshold of each ramp in isolation (leaving the others unchanged), and evaluate the achieved accuracy and latency savings as described above. Apparate then chooses the single ramp threshold change that delivered the largest additional latency savings per unit of additional accuracy loss. This process repeats until no ramp’s threshold can be increased without an accuracy violation.
@@
To enhance this process, Apparate follows a multiplicative increase, multiplicative decrease policy on step sizes to balance search speed and granularity. Each time a step increase results in an accuracy violation, Apparate halves that ramp’s step size for subsequent rounds to hone in on the boundary at fine granularity; step sizes are lower-bounded at 0.01. Conversely, selection of a ramp for threshold alteration suggests a promising path of exploration; in this case, for a speedup, Apparate doubles that ramp’s step size for the following round.
@@
Overall, as shown in Figure 10, Apparate’s threshold tuning algorithm runs up to 3 orders of magnitude faster than a pure grid search (11.9ms vs. 3.0s on average). Note that these results maximally parallelize grid search across a 32-core machine. Further, selected threshold values achieve within 0–3.8% of the latency savings of the optimal configurations.
@@
Figure 10. Apparate’s tuning vs. optimal tuning on runtime and latency of selected configurations. Bars list medians across all model-workload pairs, with error bars for min-max.
3.3. Latency-Focused Ramp Adjustments
The set of active ramps ultimately dictates where inputs can exit, and thus provides bounds on potential latency savings. Unlike threshold tuning which runs reactively (since accuracy is a constraint) and uses only recent profiling data to evaluate new configurations, ramp adjustment is used strictly as an optimization (for latency savings) in Apparate, and requires deployment to evaluate the impact of any new ramp. Thus, Apparate’s ramp tuning runs periodically (every 128 samples by default) and conservatively alters the set of active ramps to incrementally converge on high-performing configurations.
Evaluating active ramps. In each round, Apparate’s controller starts by computing a utility score for each active ramp that evaluates its overall impact on workload latency. To do so, Apparate couples per-ramp exit rates (§3.2) with two additional inputs that are collected once per model during bootstrapping: (1) the latency overhead per ramp, and (2) a layer-wise breakdown of time spent during model inference (for different batch sizes (Gujarati et al., 2020) and including network delays for distributed serving). The latter is necessary since latency characteristics vary across models but govern the impact of exits, e.g., latency arises early in CV models (Padmanabhan et al., 2022), but more evenly across coding blocks in transformers. Using these inputs, Apparate defines the utility of ramp $R$ as savings - overheads, where savings is the sum of raw latency that exiting inputs avoided by using ramp $R$, and overheads is the sum of latency that $R$ added to inputs that it could not exit.